### v1

@@ 2.2 Optimal model, barycentre model, and average drift model
(5) $\varsigma(t,x):=\tfrac{1}{2}\sum_{k\in{\mathcal{K}}}\pi_{k}\,\left(\mu^{(k)}(t,x)-{\bar{\mu}}(t,x)\right)^{\intercal}\,\Sigma^{-1}(t,x)\,\left(\mu^{(k)}(t,x)-{\bar{\mu}}(t,x)\right).$
The next result characterises the value function and the drift of the probability measure which attains the value function.
Proposition 2.2 (Value function and optimal drift).
Let Assumptions 2.1, 2.2 and 2.3 be satisfied. Then, for fixed $\eta\in{\mathds{R}}^{2}$, a candidate for the value function is
(6) $L_{\eta}(t,x)=-\log{\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}_{t,x}\left[\;e^{-\int_{t}^{T}(\varsigma(u,X_{u})\,+\eta_{0}\,g_{u})\,{\rm d}u-\eta_{1}\,f_{T}}\;\right],$
where ${\bar{\mu}}(t,x)$ and $\varsigma(t,x)$ are given in Equations (4) and (5), respectively. Moreover, a probability measure that attains the candidate value function in (3) is given by ${\mathbb{Q}}[\theta_{\eta}]$, where
(7) $\theta_{\eta}(t,x)={\bar{\mu}}(t,x)-\Sigma(t,x)\nabla L_{\eta}(t,x)\,.$
Proof 2.3.
The Hamilton-Jacobi-Bellman equation associated with the value function is
(8) $\left\{\begin{split}\partial_{t}L+\min_{\theta\in{\mathds{R}}^{d}}\left\{\theta^{\intercal}\,\nabla L+\tfrac{1}{2}{{\rm Tr}}\left(\Sigma\,\nabla^{2}L\right)+\tfrac{1}{2}\sum_{k\in{\mathcal{K}}}\pi_{k}(\theta-\mu^{(k)})^{\intercal}\Sigma^{-1}(\theta-\mu^{(k)})+\eta_{0}\,g\right\}&=0,\\ L(T,x)&=\eta_{1}\,f(x),\end{split}\right.$
where we suppressed the arguments $t,x$ and $\eta$. The unique optimiser is
$\theta^{*}={\bar{\mu}}-\Sigma\,\nabla L\,.$
Upon substituting into the inner minimisation of (8), we obtain
$\min_{\theta\in{\mathds{R}}}$$\left\{\theta^{\intercal}\,\nabla L+\tfrac{1}{2}{{\rm Tr}}\left(\Sigma\,\nabla^{2}L\right)+\tfrac{1}{2}\sum_{k\in{\mathcal{K}}}\pi_{k}(\theta-\mu^{(k)})^{\intercal}\Sigma^{-1}(\theta-\mu^{(k)})+\eta_{0}\,g\right\}$
$\begin{split}&=({\bar{\mu}}-\Sigma\,\nabla L)^{\intercal}\,\nabla L+\tfrac{1}{2}{{\rm Tr}}\left(\Sigma\,\nabla^{2}L\right)\\ &\quad+\tfrac{1}{2}\sum_{k\in{\mathcal{K}}}\pi_{k}({\bar{\mu}}-\mu^{(k)}-\Sigma\,\nabla L)^{\intercal}\Sigma^{-1}({\bar{\mu}}-\mu^{(k)}-\Sigma\,\nabla L)+\eta_{0}\,g\end{split}$
$\begin{split}&=-\nabla L^{\intercal}\,\Sigma\,\nabla L+\tfrac{1}{2}{{\rm Tr}}\left(\Sigma\,\nabla^{2}L\right)+{\bar{\mu}}^{\intercal}\nabla L\\ &\quad+\tfrac{1}{2}\sum_{k\in{\mathcal{K}}}\pi_{k}\left\{({\bar{\mu}}-\mu^{(k)})^{\intercal}\Sigma^{-1}({\bar{\mu}}-\mu^{(k)})+\nabla L^{\intercal}\,\Sigma\,\nabla L-2({\bar{\mu}}-\mu^{(k)})^{\intercal}\nabla L\right\}+\eta_{0}\,g\,.\end{split}$
Thus, the Dynamic Programming Equation becomes
$\partial_{t}L+{\bar{\mu}}^{\intercal}\,\nabla L+\tfrac{1}{2}{{\rm Tr}}\left(\Sigma\,\nabla^{2}L\right)-\tfrac{1}{2}\nabla L^{\intercal}\Sigma\nabla L+\tfrac{1}{2}\sum_{k\in{\mathcal{K}}}\pi_{k}\,\Delta\mu^{(k)^{\intercal}}\,\Sigma^{-1}\,\Delta\mu^{(k)}+\eta_{0}\,g=0,$
where we set $\Delta\mu^{(k)}:=\mu^{(k)}-{\bar{\mu}}$.
Writing $L(t,x)=-\log\omega(t,x)$, we find that $\omega$ satisfies the PDE
$\left\{\begin{split}\partial_{t}\omega+{\bar{\mu}}^{\intercal}\,\nabla\omega+\tfrac{1}{2}{{\rm Tr}}\left(\Sigma\,\nabla^{2}\omega\right)-(\varsigma+\eta_{0}\,g)\,\omega&=0,\\ \omega(T,x)&=e^{-\eta_{1}\,f(x)}\,.\end{split}\right.$
The Feyman-Kac formula then implies that we may represent
$\omega(t,x)={\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}_{t,x}\left[e^{-\int_{t}^{T}(\varsigma(u,X_{u})+\eta_{0}\,g_{u})\,{\rm d}u-\eta_{1}\,f(X_{T})}\right]\,,$
which concludes the proof.
We provide a verification result shortly.
To obtain the solution to problem (P), the optimal Lagrange multipliers $\eta^{*}$ must be chosen to bind the constraints, that is solve
${\mathbb{E}}^{{\mathbb{Q}}[\theta_{\eta^{*}}]}\Big{[}\int_{0}^{T}g_{u}\,{\rm d}u\Big{]}=0\qquad\text{and}\qquad{\mathbb{E}}^{{\mathbb{Q}}[\theta_{\eta^{*}}]}\big{[}f_{T}\big{]}=0\,\,.$
Before stating the solution to optimisation problem (P), we consider the (pure) barycentre measure without any constraints, which is essential in determining the optimal Lagrange parameters.
Definition 2.4 (Barycentre measure).
Consider optimisation problem (P) without any constraints, whose solution we call the barycentre (probability) measure and denote it by ${\mathbb{Q}}^{\mathfrak{b}}$, that is
The Radon-Nikodym (RN) derivative of the barycentre model to the average drift model admits a succinct representation as presented next.
Proposition 2.5 (Pure barycentre).
Let Assumptions 2.1, 2.2 and 2.3 be satisfied. Then the barycentre measure exists, is unique, and its RN derivative has representation
@@
(9) $\frac{{\rm d}{\mathbb{Q}}^{\mathfrak{b}}}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}=\frac{e^{-\int_{0}^{T}\varsigma(t,X_{t})\,{\rm d}t}}{{\mathbb{E}}^{{}^{{\mathbb{Q}}[{\bar{\mu}}]}}\left[e^{-\int_{0}^{T}\varsigma(t,X_{t})\,{\rm d}t}\right]}\;.$
Proof 2.6.
Existence and uniqueness follows by the strict convexity and coerciveness of the KL divergence. Setting $\eta_{0}=\eta_{1}=0$ in (3) corresponds to solving for the barycentre measure. From Proposition 2.2 and we have that the Lagrangian $L_{0}(t,x)=-\log{\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}_{t,x}\left[\;e^{-\int_{t}^{T}\varsigma(u,X_{u})\,{\rm d}u}\;\right]$ and by 2.3, $L_{0}(t,x)$ is finite for all $t$ and $x$. The representation (9) follows by Theorem 2.7 in [15].
We remark that the RN derivative of the barycentre to the average drift model is of Escher type, but different to the solution to the classical entropy maximisation, see e.g., [9]. Interestingly, in the exponent is a time integral over the weighted deviation between each expert’s drift to the average drift ${\bar{\mu}}(t,x)$ scaled by the instantaneous covariance matrix $\Sigma(t,x)$.
@@
Next, we make the following assumption, which guarantees that the constraints imposed by the agent are feasible.
@@
We denote the solution by $\eta^{*}:=(\eta^{*}_{0},\eta^{*}_{1})$.
Theorem 2.7 (Verification).
Let Assumptions 2.1, 2.2, and 2.3 be satisfied. Take $\eta\in{\mathds{R}}^{2}$ such that ${\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[\exp\{-\eta_{0}\int_{0}^{T}g_{t}\,{\rm d}t-\eta_{1}\,f_{T}\}]<+\infty$. Suppose that $L_{\eta}(t,x)$ in (6) is in ${\mathcal{C}}^{1,2}([0,T)\times{\mathds{R}}^{d};{\mathds{R}})\cap{\mathcal{C}}^{0}([0,T]\times{\mathds{R}}^{d};{\mathds{R}})$ and has at most quadratic growth, i.e. $|L_{\eta}(t,x)|\leq C(1+|x|^{2})$ for some $C\in{\mathds{R}}_{+}$. Let $\theta_{\eta}(t,x)$ be as in (7), and suppose that
${\mathbb{E}}^{{\mathbb{Q}}[\theta_{\eta}]}\left[e^{\frac{1}{2}\int_{0}^{T}|{\bar{\lambda}}_{u}|^{2}{\rm d}u}\right]<+\infty\,,\quad{\mathbb{E}}^{{\mathbb{Q}}[\theta_{\eta}]}\left[\sup_{t\in[0,T]}|X_{t}|^{2}\right]<+\infty\,,$
@@
where ${\bar{\lambda}}(t,x):=\sigma^{-1}(t,x)\big{(}{\bar{\mu}}(t,x)-\theta_{\eta}(t,x)\big{)}$. Then $\theta_{\eta}(t,x)$ is admissible and $L_{\eta}(t,x)$ is the value function.
Proof 2.8.
The proof follows along the lines of Theorem 2.6 in [15], and is omitted.
We are finally in the position to state the succinct representation of the change of measure of the optimal measure to the average drift model.
@@
Let Assumptions 2.1, 2.2, 2.3, and 2.4 be satisfied. Then, there exists a unique solution to optimisation problem (P). The measure that attains the infimum has RN derivative given by
(11) $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}=\frac{e^{-\int_{0}^{T}(\varsigma(t,X_{t})+\eta^{*}_{0}\,g_{t})\,{\rm d}t-\eta_{1}^{*}\,f_{T})}}{{\mathbb{E}}^{{}^{{\mathbb{Q}}[{\bar{\mu}}]}}\left[e^{-\int_{0}^{T}(\varsigma(t,X_{t})+\eta_{0}^{*}\,g_{t})\,{\rm d}t-\eta_{1}^{*}\,f_{T}}\right]}\;.$
Proof 2.10.
For any $\eta$ the solution to (3) is ${\mathbb{Q}}[\theta_{\eta}]$ and given by Proposition 2.2. Thus, we only need to find the Lagrange multipliers $\eta$ that bind the constraints, which then gives the solution to optimisation problem (P). Uniqueness follows by strict convexity of the KL divergence and as the constraints are linear functionals of the probability measure.
To obtain an explicit representation of $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$, we use the optimal drift representation in Proposition 2.2 and apply Theorem 2.7 in [15], to obtain that
@@
Next, using the above representation of the RN derivative, we rewrite the running constraint expectation

### v2

@@ 3.2 Ornstein-Uhlenbeck experts’ processes
$\zeta_{t}:=\frac{1}{2\sigma^{2}_{t}}\sum_{k\in{\mathcal{K}}}\pi_{k}\,\big{(}\beta_{t}^{(k)}\big{)}^{2}.$
Let $\tilde{{\mathfrak{c}}_{t}}:=e^{\alpha_{t}}\,{\mathfrak{c}}_{t}$, then we have
$\tfrac{{\rm d}}{{\rm d}t}(\tilde{\mathfrak{c}}_{t})=-\tfrac{1}{2}e^{-\alpha_{t}}\sigma^{2}_{t}\,\tilde{\mathfrak{c}}_{t}^{2}+e^{\alpha_{t}}\,\zeta_{t}$
Defining, $\tilde{\sigma}_{t}:=\tfrac{1}{2}e^{-\alpha_{t}}\sigma^{2}_{t}$ and $\tilde{\zeta}_{t}:=e^{\alpha_{t}}\,\zeta_{t}$, we may re-write the above as
$\tfrac{{\rm d}}{{\rm d}t}(\tilde{\mathfrak{c}}_{t})=-\tilde{\sigma}_{t}\,\tilde{\mathfrak{c}}_{t}^{2}+\tilde{\zeta}_{t}.$
Note, the coefficient of $\tilde{\mathfrak{c}}_{t}^{2}$ on the rhs is negative, consequently,
(54) $\tfrac{{\rm d}}{{\rm d}t}(\tilde{\mathfrak{c}}_{t})\leq\tilde{\zeta}_{t}\qquad\Rightarrow\qquad\tilde{\mathfrak{c}}_{T}-\tilde{\mathfrak{c}}_{t}\leq\int_{t}^{T}\tilde{\zeta}_{s}\,{\rm d}s\qquad\Rightarrow\qquad\tilde{\mathfrak{c}}_{t}\geq-\int_{t}^{T}\tilde{\zeta}_{s}\,{\rm d}s$
This provides a lower bound on $\tilde{{\mathfrak{c}}}_{t}$ and hence a lower bound of ${\mathfrak{c}}_{t}$. Next, we obtain an upper bound. By (51), we deduce that $\left.\left(\frac{{\rm d}}{{\rm d}t}\tilde{\mathfrak{c}}_{t}\right)\right|_{t=T}=\zeta_{T}>0$, hence there exists $\varepsilon>0$, s.t. $\forall t\in[T-\varepsilon,T)$, $\tilde{\mathfrak{c}}_{t}<0$.
Next, take $s=\sup\{u\in[0,T):\tilde{\mathfrak{c}}_{u}=0\}$. Then, by (51), we have $\left.\left(\frac{{\rm d}}{{\rm d}t}\tilde{\mathfrak{c}}_{t}\right)\right|_{t=s}=\zeta_{s}>0$, and hence there exists $\varepsilon_{s}>0$, s.t. $\forall u\in[s-\varepsilon_{s},s)$, $\tilde{\mathfrak{c}}_{u}<0$. Repeating this argument for all times $(s_{i})_{i\in\mathbb{N}}$ s.t. $\tilde{\mathfrak{c}}_{s_{i}}$ is zero, we conclude that $\tilde{\mathfrak{c}}_{t}\leq 0$ for all $t\in[0,T]$. Combining with the inequality (54), we have
$-\int_{t}^{T}\tilde{\zeta}_{s}\,{\rm d}s\leq\tilde{\mathfrak{c}}_{t}\leq 0,$
and therefore we arrive at the bound
(55) $-e^{-\alpha_{t}}\int_{t}^{T}\tilde{\zeta}_{s}\,{\rm d}s\leq{\mathfrak{c}}_{t}\leq 0.$
As ${\mathfrak{c}}_{t}$ remains finite on the interval $t\in[0,T]$, we may solve for ${\mathfrak{b}}_{t}$ by using an integrating factor as follows:
(56) $\frac{d}{dt}\left(e^{-\int_{t}^{T}(\sigma^{2}_{u}\,{\mathfrak{c}}_{u}-{\overline{b}}_{u})\,{\rm d}u}\,{\mathfrak{b}}_{t}\right)$ $=\left((\sigma^{2}_{t}\,{\mathfrak{c}}_{t}-{\overline{b}}_{t})\,\,{\mathfrak{b}}_{t}+\dot{\mathfrak{b}}_{t}\right)e^{-\int_{t}^{T}(\sigma^{2}_{u}{\mathfrak{c}}_{u}-{\overline{b}}_{u})\,{\rm d}u}$
(57) $=-\left({\overline{a}}_{t}\,{\mathfrak{c}}_{t}+\frac{1}{\sigma^{2}_{t}}\sum_{k\in{\mathcal{K}}}\pi_{k}\,\alpha_{t}^{(k)}\,\beta_{t}^{(k)}\right)e^{-\int_{t}^{T}(\sigma^{2}_{u}{\mathfrak{c}}_{u}-{\overline{b}}_{u})\,{\rm d}u}\,.$
Integrating from $t$ to $T$, and as ${\mathfrak{b}}_{T}=0$, we obtain
(58) $-e^{-\int_{t}^{T}(\sigma^{2}_{u}\,{\mathfrak{c}}_{u}-{\overline{b}}_{u})\,{\rm d}u}\,{\mathfrak{b}}_{t}=-\int_{t}^{T}\left({\overline{a}}_{s}\,{\mathfrak{c}}_{s}+\frac{1}{\sigma^{2}_{s}}\sum_{k\in{\mathcal{K}}}\pi_{k}\,\alpha_{s}^{(k)}\,\beta_{s}^{(k)}\right)e^{-\int_{s}^{T}(\sigma^{2}_{u}{\mathfrak{c}}_{u}-{\overline{b}}_{u})\,{\rm d}u}\,{\rm d}s$
which implies
(59) ${\mathfrak{b}}_{t}=\int_{t}^{T}\left({\overline{a}}_{s}\,{\mathfrak{c}}_{s}+\frac{1}{\sigma^{2}_{s}}\sum_{k\in{\mathcal{K}}}\pi_{k}\,\alpha_{s}^{(k)}\,\beta_{s}^{(k)}\right)e^{\int_{t}^{s}(\sigma^{2}_{u}{\mathfrak{c}}_{u}-{\overline{b}}_{u})\,{\rm d}u}\,{\rm d}s.$
Then for ${\mathfrak{a}}_{t}$, integrating from $t$ to $T$, and as ${\mathfrak{a}}_{T}=0$, we obtain
@@
(60) ${\mathfrak{a}}_{t}=\int_{t}^{T}\left({\overline{a}}_{s}\,{\mathfrak{b}}_{s}+\tfrac{1}{2}\sigma^{2}_{s}({\mathfrak{c}}_{s}+{\mathfrak{b}}_{s}^{2})-\frac{1}{2\sigma^{2}_{s}}\sum_{k\in{\mathcal{K}}}\pi_{k}\,\big{(}\alpha^{(k)}_{s}\big{)}^{2}\right)\,{\rm d}s\,.$
Due to the boundedness on the coefficients $(a^{(k)},b^{(k)})_{k\in{\mathcal{K}}}$ and the established bound on ${\mathfrak{c}}$, we see that ${\mathfrak{a}}$ and ${\mathfrak{b}}$ are also bounded on the interval $[0,T]$.
Therefore, $L_{0}(t,x)=-({\mathfrak{a}}_{t}+{\mathfrak{b}}_{t}\,x+\frac{1}{2}{\mathfrak{c}}_{t}\,x^{2})$ and from (9) we obtain the stated result.
The next proposition establishes that this class of OU expert models satisfy the requirements of Theorem 3.2.
Proposition 3.12.
For the expert models specified above, we have that $L_{0}\in{\mathcal{C}}^{1,2}([0,T)\times{\mathds{R}}^{d};{\mathds{R}})\cap{\mathcal{C}}^{0}([0,T]\times{\mathds{R}}^{d};{\mathds{R}})$ and has at most quadratic growth, and
@@
${\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[e^{\frac{1}{2}\int_{0}^{T}|{\bar{\lambda}}_{0,u}|^{2}{\rm d}u}\right]<+\infty\,,\quad{\mathbb{E}}^{{\mathbb{Q}}[\theta_{0}]}\left[\sup_{t\in[0,T]}|X_{t}|^{2}\right]<+\infty\,,$
where ${\bar{\lambda}}_{0,t}:=\sigma^{-1}_{t}\big{(}{\bar{\mu}}_{t}-\theta_{0,t}\big{)}$ and
(61) $\theta_{0,t}:={\bar{\mu}}_{t}-\Sigma_{t}\nabla_{x}L_{0}(t,X_{t})\,.$
Proof 3.13.
First, form Proposition 3.10, we see that $L_{0}$ satisfies the properties in the proposition statement. Next, there exists $C$, s.t.,
(62) ${\mathbb{E}}^{{\mathbb{Q}}[\theta_{0}]}\left[\sup_{t\in[0,T]}|X_{t}|^{2}\right]\leq C\,{\mathbb{E}}^{{\mathbb{Q}}[\theta_{0}]}\left[\int_{0}^{T}d[X,X]_{t}\right]=C\,{\mathbb{E}}^{{\mathbb{Q}}[\theta_{0}]}\left[\int_{0}^{T}\sigma_{t}^{2}\,{\rm d}t\right]<\infty\,.$
Next, by Proposition 3.10, ${\bar{\lambda}}_{0,t}=\sigma_{t}\nabla_{x}L_{0}(t,X_{t})=-\sigma_{t}({\mathfrak{b}}_{t}+{\mathfrak{c}}_{t}\,X_{t})$, therefore,
(63) ${\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[e^{\frac{1}{2}\int_{0}^{T}|{\bar{\lambda}}_{0,u}|^{2}{\rm d}u}\right]$ $={\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[e^{\frac{1}{2}\int_{0}^{T}\sigma_{u}^{2}({\mathfrak{b}}_{u}+{\mathfrak{c}}_{u}\,X_{u})^{2}{\rm d}u}\right]\,.$
Define $h(t,x):={\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}_{t,x}\left[e^{\frac{1}{2}\int_{t}^{T}\sigma_{u}^{2}({\mathfrak{b}}_{u}+{\mathfrak{c}}_{u}\,X_{u})^{2}{\rm d}u}\right]$, then $h$ satisfies the PDE
(64) $\left\{\begin{array}[]{rr}\partial_{t}h+{\bar{\mu}}\,\partial_{x}h+\tfrac{1}{2}\sigma^{2}\,\partial_{xx}h+\frac{1}{2}\sigma^{2}({\mathfrak{b}}+{\mathfrak{c}}x)^{2}&=0\,,\\ h(T,x)&=1\,.\end{array}\right.$
As this is an affine PDE, it admits a solution of the form $h(t,x)=e^{{\mathfrak{l}}_{t}+{\mathfrak{f}}_{t}\,x+{\mathfrak{g}}_{t}\,x^{2}}$, where ${\mathfrak{l}}$, ${\mathfrak{f}}$, and ${\mathfrak{g}}$ are all deterministic functions of time. As ${\bar{\mu}}$, $\sigma$, ${\mathfrak{b}}$, and ${\mathfrak{c}}$ are all bounded, using similar arguments as in the proof of Proposition 3.10, we find that ${\mathfrak{l}}$, ${\mathfrak{f}}$, and ${\mathfrak{g}}$ are all bounded for all $t\in[0,T]$. Therefore, we conclude ${\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[e^{\frac{1}{2}\int_{0}^{T}|{\bar{\lambda}}_{0,u}|^{2}{\rm d}u}\right]=h(T,x)<\infty$.
4 The case with beliefs
In this section, we extend the barycentre discussed above to the case when the agent has additional beliefs that they wish to incorporate. That is, we solve (P).
4.1 Optimal model
@@
Beliefs may be incompatible with the experts’ models. To avoid this scenario, we make the following assumption, which guarantees that the constraints imposed by the agent are feasible.
@@
Lemma 4.1.
If the random variable $Y:=(\int_{0}^{T}g_{t}\,{\rm d}t,f_{T})^{\intercal}$ is ${\mathbb{Q}}^{\mathfrak{b}}$–non-degenerate:
$\forall c\in{\mathds{R}}^{2},\qquad{\mathbb{Q}}^{\mathfrak{b}}\left(\{\omega:Y(\omega)=c\}\right)<1,$
i.e., $Y$ is ${\mathbb{Q}}^{\mathfrak{b}}$–a.s. not constant, then if there exists a solution to (65), it is unique and we denote this unique solution by $\eta^{*}:=(\eta^{*}_{0},\eta^{*}_{1})$.
Proof 4.2.
Let $K(\eta):={\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[e^{-\eta\cdot Y}]$. We first show that $K(\eta)$ is strictly convex. To this end, note that
(66) $\nabla_{\eta}K(\eta)=-\frac{{\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[Y\,e^{-\eta\cdot Y}]}{{\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[e^{-\eta\cdot Y}]}=-{\mathbb{E}}^{{{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}}\left[Y\right]$
and
$\nabla^{2}_{\eta\eta}K(\eta)$ $=\frac{{\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[Y\,Y^{\intercal}\,e^{-\eta\cdot Y}]}{{\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[e^{-\eta\cdot Y}]}-\frac{{\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[Y\,e^{-\eta\cdot Y}]{\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[Y^{\intercal}\,e^{-\eta\cdot Y}]}{\left({\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[e^{-\eta\cdot Y}]\right)^{2}}$
(67) $={\mathbb{E}}^{{{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}}[Y\,Y^{\intercal}]-{\mathbb{E}}^{{{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}}[Y]\,{\mathbb{E}}^{{{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}}[Y^{\intercal}]$
(68) $={\rm{Cov}}^{{{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}}[Y]$
where the measure ${{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}$ is defined by the Radon-Nikodym derivative
$\frac{{\rm d}{\mathbb{Q}}_{\eta}}{{\rm d}{\mathbb{Q}}^{\mathfrak{b}}}=\frac{e^{-\eta\cdot Y}}{{\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[e^{-\eta\cdot Y}]}\,,$
and ${\rm{Cov}}^{{{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}}[Y]$ denotes the ${{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}$-covariance matrix of $Y$. As $Y$ is ${{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}$–non-degenerate and as ${{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}$ and ${\mathbb{Q}}^{\mathfrak{b}}$ are equivalent, it is also ${\mathbb{Q}}^{\mathfrak{b}}_{\eta}$–non-degenerate. Hence, ${\rm{Cov}}^{{{\mathbb{Q}}^{\mathfrak{b}}_{\eta}}}[Y]$ and therefore $\nabla_{\eta\eta}^{2}K(\eta)$, is strictly positive definite. Therefore $K(\eta)$ is strictly convex. Consequently, if a minimiser of $K$ exists, it is unique, and solves the first order condition $\nabla_{\eta}K(\eta)=0$. Thus, there can be at most one solution to (65).
The theorem below provides our next key result the solution to optimisation problem (P), i.e., the drift under the optimal measure, given in (70), and the associated value function, given in (69).
Theorem 4.3 (Optimal Drift and Value Function with Beliefs).
Let Assumptions 2.1 and 2.2 be satisfied. Take $\eta\in{\mathds{R}}^{2}$ such that ${\mathbb{E}}^{{\mathbb{Q}}^{\mathfrak{b}}}[\exp\{-\eta_{0}\int_{0}^{T}g_{t}\,{\rm d}t-\eta_{1}\,f_{T}\}]<+\infty$.
Define the function $L_{\eta}:[0,T]\times{\mathds{R}}^{d}\to{\mathds{R}}$, s.t.
(69) $L_{\eta}(t,x):=-\log{\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}_{t,x}\left[\;e^{-\int_{t}^{T}(\varsigma(u,X_{u})\,+\eta_{0}\,g_{u})\,{\rm d}u-\eta_{1}\,f_{T}}\;\right],\qquad\forall(t,x)\in[0,T]\times{\mathds{R}}^{d}\,.$
Suppose that $L_{\eta}\in{\mathcal{C}}^{1,2}([0,T)\times{\mathds{R}}^{d};{\mathds{R}})\cap{\mathcal{C}}^{0}([0,T]\times{\mathds{R}}^{d};{\mathds{R}})$ and has at most quadratic growth, i.e. there exists $C\in{\mathds{R}}_{+}$ s.t. $|L_{\eta}(t,x)|\leq C(1+|x|^{2})$ for all $(t,x)\in[0,T]\times{\mathds{R}}^{d}$. Next, define the process $\theta_{\eta}:=(\theta_{\eta,t})_{{t\in[0,T]}}$
(70) $\theta_{\eta,t}:={\bar{\mu}}_{t}-\Sigma_{t}\nabla_{x}L_{\eta}(t,X_{t})\,,$
and suppose that
${\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[e^{\frac{1}{2}\int_{0}^{T}|{\bar{\lambda}}_{\eta,u}|^{2}{\rm d}u}\right]<+\infty\,,\quad{\mathbb{E}}^{{\mathbb{Q}}[\theta_{\eta}]}\left[\sup_{t\in[0,T]}|X_{t}|^{2}\right]<+\infty\,,$
@@
where ${\bar{\lambda}}_{\eta,t}:=\sigma^{-1}_{t}\big{(}{\bar{\mu}}_{t}-\theta_{\eta,t}\big{)}$. Then $\theta_{\eta}$ is admissible, $L_{\eta}$ is the value function in (6), and the probability measure that attains the minimum is ${\mathbb{Q}}[\theta_{\eta}]$.
Proof 4.4.
The proof follows along the lines of Theorem 3.2 and is omitted for brevity.
Similarly to the pure barycentre case, we have the following representation for the RN derivative of the optimal measure.
Proposition 4.5 (Measure change representation).
Under the assumptions of Theorem 4.3, and for $\eta\in{\mathds{R}}^{2}$, the RN derivative $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}$ has representation
(71) $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}=\frac{e^{-\int_{0}^{T}(\varsigma(t,X_{t})+\eta_{0}\,g_{t}\,{\rm d}t-\eta_{1}\,f_{T}}}{{\mathbb{E}}^{{\mathbb{Q}}[{\bar{\mu}}]}\left[e^{-\int_{0}^{T}(\varsigma(t,X_{t})+\eta_{0}\,g_{t}\,{\rm d}t-\eta_{1}\,f_{T}}\right]}\;.$
Proof 4.6.
The proof follows along the same lines as Proposition 3.4, and is omitted for brevity.
Thus far, we have the value function, drift, and optimal measure identified for problem (6) for an arbitrary $\eta$. To obtain the solution to orignal problem (P), the optimal Lagrange multipliers $\eta^{\star}$ must be chosen to bind the constraints, that is to solve
${\mathbb{E}}^{{\mathbb{Q}}[\theta_{\eta^{\star}}]}\Big{[}\int_{0}^{T}g_{u}\,{\rm d}u\Big{]}=0\qquad\text{and}\qquad{\mathbb{E}}^{{\mathbb{Q}}[\theta_{\eta^{\star}}]}\big{[}f_{T}\big{]}=0\,\,.$
The following result shows that the probability measure that solves (P) exists and is unique. Moreover, it has a RN derivative, relative to the average drift measure, of the exponential type, similar to the barycentre measure.
@@
Let the assumptions in Theorem 4.3 and 4.1 hold. Then, there exists a unique solution to optimisation problem (P). Moreover, the measure that attains the infimum has RN derivative given by
(72) $\frac{{\rm d}{\mathbb{Q}}[\theta_{\eta^{*}}]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}=\frac{e^{-\int_{0}^{T}(\varsigma(t,X_{t})+\eta^{*}_{0}\,g_{t})\,{\rm d}t-\eta_{1}^{*}\,f_{T})}}{{\mathbb{E}}^{{}^{{\mathbb{Q}}[{\bar{\mu}}]}}\left[e^{-\int_{0}^{T}(\varsigma(t,X_{t})+\eta_{0}^{*}\,g_{t})\,{\rm d}t-\eta_{1}^{*}\,f_{T}}\right]}\;,$
where $\eta^{*}$ solves (65) in 4.1.
Proof 4.8.
For any $\eta$ the solution to (6) is ${\mathbb{Q}}[\theta_{\eta}]$ and given by (71). Thus, we only need to find the Lagrange multipliers $\eta$ that bind the constraints, which then gives the solution to optimisation problem (P). Uniqueness follows by strict convexity of the KL divergence and as the constraints are linear functionals of the probability measure.
For $\eta$ fixed, Proposition 4.5 states that
@@
Continuing, using the above representation of the RN derivative, we rewrite the running constraint